OpenStudy (anonymous):

1/2(x+3)-2=1 what is the solution

OpenStudy (anonymous):

please help keep getting the wrong answer

OpenStudy (luigi0210):

Write ot using the equaion editor

OpenStudy (luigi0210):

*equation

OpenStudy (ja1):

\[\frac{ 1 }{ 2 }(x+3)-2=1\]

OpenStudy (ja1):

Is that it @jonny0905 ?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

Well then simple, just multiply by 2 to get rid of the fraction and solve from there.

OpenStudy (anonymous):

they say the answer is x = 3

OpenStudy (luigi0210):

Did you solve it out?

OpenStudy (anonymous):

yes its still not right...i think. i did this 2/1(1/2)= 1 1(x+3)-2=1 x + 1 = 1 x = 0 than that didnt work so i did this 2(1/2)((2)x+2(3))-2(2)=2(1) 2x + 6 - 4 = 2 2x + 2 = 2 2x = 0 its supposed to be x = 3

OpenStudy (luigi0210):

I see where you messed up, you didn't multiply both sides

OpenStudy (anonymous):

by what? 2?

OpenStudy (luigi0210):

Yup

OpenStudy (anonymous):

i did in the second one i multiplied everything by 2

OpenStudy (anonymous):

can you show your steps?

OpenStudy (anonymous):

idk why this is giving me trouble its so simple

OpenStudy (luigi0210):

 Finish it?